OpenStudy (newwar):

At 3:00 p.m., Jamie leaves her house and drives north to the beach. She arrives at 5:30 p.m. Also leaving at 3:00 p.m., her brother Raul leaves her house and drives south to a state park. He arrives at 4:30 p.m. The driving distance from the beach to the state park is 240 mi. Jamie and Raul drove at the same speed. What was their speed? Enter your answer in the box. mph

OpenStudy (newwar):

@ShadowLegendX

OpenStudy (newwar):

@Goldendragon01

OpenStudy (newwar):

@Helpmeeasetjx

OpenStudy (jiteshmeghwal9):



OpenStudy (jiteshmeghwal9):

Let the speed of both Jamie & her brother be 'v' miles per hour. assume distance between Jamie's house & beach =x miles then distance between Jamie's house & state park= (240-x) miles Time taken by Jamie to reach the beach = 2.5 hour Time taken by Jamie's brother to reach state park = 1.5 hour

OpenStudy (jiteshmeghwal9):

\[\text{Speed}=\frac{\text{distance traveled}}{\text{time taken }}\] distance traveled by Jamie in 2.5 hour at a constant speed 'v' miles per hour is x miles. so, v=\(\Large \frac{x}{2.5}\).......(1) The distance traveled by Jamie's brother in 1.5 hour at v a constant speed 'v' is (240-x) miles v=\(\Large \frac{240-x}{1.5}\) ....(2)

OpenStudy (jiteshmeghwal9):

solve equation (1) & (2) for 'v'.